# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ Abstract
Recently, Kleshchev and Livesey [21, 22] proved the existence of RoCK $p$-blocks for double covers of symmetric and alternating groups over large enough coefficient rings. They proved that RoCK blocks of double covers are Morita equivalent to standard “local” blocks via bimodules with endopermutation source. Based on this, Kleshchev and Livesey proved that these RoCK blocks are splendidly Rickard equivalent to their Brauer correspondents. The analogous result for blocks of symmetric groups, a theorem of Chuang and Kessar, was an important step in Chuang and Rouquier ultimately proving Broué’s abelian defect group conjecture for symmetric groups. In this paper we show that in most cases the Morita equivalences and splendid Rickard equivalences constructed by Kleshchev and Livesey descend to the ring $\mathbb{Z}_{p}$ of $p$-adic integers, hence prove Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture for most of these RoCK blocks.
@@ 1 Introduction
Comparing to the refined version, the usual Broué conjecture is with the assumption that the complete discrete valuation rings is large enough. Conjecture 1.1 has been proved for blocks with cyclic defect groups in [19], blocks with a Klein four defect group in [12], blocks of alternating groups in [13], blocks of ${\rm SL}_{n}(q)$ and ${\rm GL}_{n}(q)$ in defining characteristic (see [15]), and unipotent blocks ${\rm GL}_{n}(q)$ in non-defining characteristic (see [16]). Conjecture 1.1 has also been verified for certain blocks of $p$-nilpotent and $p$-solvable groups in [3], [27] and [28]. Since any field is a splitting field for symmetric groups, Conjecture 1.1 has been proved for blocks of symmetric groups by Chuang and Rouquier [6, Theorem 7.6].
@@
For a non-negative integer $n$, denote by $S_{n}$ the symmetric group on $n$ letters and by $\widetilde{S}_{n}$ one of the two double covers $\widetilde{S}_{n}^{+}$ and $\widetilde{S}_{n}^{-}$ of $S_{n}$. Denote by $A_{n}$ the corresponding alternating group and by $\widetilde{A}_{n}$ its double cover. In this paper we provide additional evidence for Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture by considering blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$.
@@
Let $\mathcal{O}$ be a complete discrete valuation ring. Denote by $z$ the central element in $\widetilde{S}_{n}$ of order $2$. Since $\mathcal{O}\widetilde{S}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}S_{n}$ and $\mathcal{O}\widetilde{A}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}A_{n}$, Conjecture 1.1 is true for blocks of $\mathcal{O}\widetilde{S}_{n}(1+z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1+z)/2$. So we restrict our attention in this paper to blocks of $\mathcal{O}\widetilde{S}_{n}(1-z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1-z)/2$, which are called the spin blocks of the symmetric and alternating groups, respectively. For block with trivial defect group, Conjecture 1.1 is trivially true, hence we only need to consider Conjecture 1.1 for block with nontrivial defect groups. Let $K$ denote the field of fractions of $\mathcal{O}$ and let $k$ denote the residue field of $\mathcal{O}$, i.e. $k=\mathcal{O}/J(\mathcal{O})$. Let $p:={\rm char}(k)$, a prime number. For the rest of this introduction we assume that $K$ has characteristic 0, and when dealing with a finite group $G$, we will assume that $K$ (and consequently $\mathcal{O}$) contains a primitive $|G|$-th root of unity; we also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$). Let $\mathbb{Z}_{p}$ denote the ring of $p$-adic integers, $\mathbb{Q}_{p}$ the field of $p$-adic numbers and $\mathbb{F}_{p}$ the field with $p$ elements. Since $z$ is a central element of order $2$, if $p=2$, then $\widetilde{S}_{n}$ (resp. $\widetilde{A}_{n}$) satisfies Conjecture 1.1 if and only if $S_{n}\cong\widetilde{S}_{n}/\langle z\rangle$ (resp. $A_{n}\cong\widetilde{A}_{n}/\langle z\rangle$) satisfies Conjecture 1.1; see [31, Lemma 10.2.11] (we note that the blanket assumption in [31] that the coefficient ring is large enough is not used in the proof of [31, Lemma 10.2.11]). Hence Conjecture 1.1 holds for $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ if $p=2$. So from now on we assume that $p\geq 3$.
@@
Let $R\in\{\mathcal{O},k\}$, $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$ and let $e_{z}:=(1-z)/2\in R\widetilde{S}_{n}$. From [21, 22] we see that it is important to consider $R\tilde{S}_{n}$ as a superalgebra via
$(R\widetilde{S}_{n})_{\bar{0}}:=R\widetilde{A}_{n}~~~{\rm and}~~~(R\widetilde{S}_{n})_{\bar{1}}:=R(\widetilde{S}_{n}\backslash\widetilde{A}_{n}).$
@@
Recall from [22] that the spin superblock algebras (i.e. the indecomposable superideals) of $R\tilde{S}_{n}e_{z}$ are labeled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core ($p$-bar core) and $d\in\mathbb{N}$ (called the weight of the block) such that $n=|\rho|+d\rho$. Denote such a superblock algebra by $B^{\rho,d}_{R}$ and denote by $e_{\rho,d}$ the central idempotent of $R\widetilde{S}_{n}$ such that $B^{\rho,d}_{R}=R\widetilde{S}_{n}e_{\rho,d}$. We prove that the idempotent $e_{\rho,d}$ is contained in the subring $R_{p}\widetilde{S}_{n}$ (see Proposition 5.9 below). Hence we can define $B^{\rho,d}_{R_{p}}:=R_{p}\widetilde{S}_{n}e_{\rho,d}$. Note that the case $d=0$ corresponds to the defect zero situations, so we assume that $d>0$ in the rest of this section. In this case $B^{\rho,d}_{R}$ is in fact a spin block algebra (and not just a superblock algebra) of $R\widetilde{S}_{n}$.
@@
We denote by $D^{\rho,d}$ a defect group of $B^{\rho,d}_{R}$. The defect group $D^{\rho,d}$ of $B^{\rho,d}_{R}$ is abelian if and only if $d<p$. Moreover, if $d<p$, then $D^{\rho,d}\cong(C_{p})^{\times d}$. By [22, Remark 4.4], the idempotent $e_{\rho,d}$ is contained in the subalgebra $R\widetilde{A}_{n}e_{z}$, and hence (by Proposition 5.9) in $R_{p}\widetilde{A}_{n}e_{z}$. By [22, Theorem 4.5] (or [18, Propsition 3.16]), as an element of $R\widetilde{A}_{n}$, $e_{\rho,d}$ is a single block unless $d=0$ and $\rho$ is even. So $B^{\rho,d}_{R}$ can be regarded as an $R$-superalgebra with even part $(B^{\rm\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}$. We denote $R_{p}\widetilde{A}_{n}e_{\rho,d}$ by $(B^{\rm\rho,d}_{R_{p}})_{\bar{0}}$. It is known that $D^{\rho,d}$ is also a defect group of $B^{\rho,d}_{\bar{0}}$ (see [18, Proposition 3.16] or Theorem 5.8 below). Let $b^{\rho,d}_{R}$ (resp. $(b^{\rho,d}_{R})_{\bar{0}}$) be the Brauer correspondent of $B^{\rho,d}_{R}$ (resp. $(B^{\rho,d}_{R})_{\bar{0}}$) in $RN_{\widetilde{S}_{n}}(D^{\rho,d})$ (resp. $RN_{\widetilde{A}_{n}}(D^{\rho,d})$). Similarly, we let $b^{\rho,d}_{R_{p}}$ (resp. $(b^{\rho,d}_{R_{p}})_{\bar{0}}$) be the Brauer correspondent of $B^{\rho,d}_{R_{p}}$ (resp. $(B^{\rho,d}_{R_{p}})_{\bar{0}}$) in $R_{p}N_{\widetilde{S}_{n}}(D^{\rho,d})$ (resp. $R_{p}N_{\widetilde{A}_{n}}(D^{\rho,d})$). Note that $(b^{\rho,d}_{R})_{\bar{0}}$ (resp. $(b^{\rho,d}_{R_{p}})_{\bar{0}}$) is indeed the even part of the superalgebra $B^{\rho,d}_{R}$ (resp. $B^{\rho,d}_{R_{p}}$). Let $\mathcal{T}_{n}^{R}:=R\widetilde{S}_{n}e_{z}$. For a superalgebra $A$, a twisted wreath superproduct $A\wr_{s}\mathcal{T}_{d}^{R}$ is defined in [22, §3.7] (see also 2.7 below).
In [21], Kleshchev and Livesey proved the following structure theorem of $b^{\rho,d}_{k}$:
Theorem 1.2 ([21, Theorem A]).
@@
    $b^{\rho,d}_{k}\cong B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$ as $k$-superalgebras.
@@
    $(b^{\rho,d}_{k})_{\bar{0}}\cong(B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}$ as $k$-algebras.
Based on the above theorem we prove a similar structure theorem of $b^{\rho,d}_{\mathbb{F}_{p}}$:
@@
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then
  1. 1.
    $b^{\rho,d}_{\mathbb{F}_{p}}\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$ as $\mathbb{F}_{p}$-superalgebras.
  2. 2.
    $(b^{\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}\cong(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$ as $\mathbb{F}_{p}$-algebras.
In [21], Kleshchev and Livesey proved the following analogue of “the Rickard and Marcus step for symmetric groups” for double covers of symmetric and alternating groups:
Theorem 1.4 ([21, Theorem C]).
@@
    $B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$ is derived equivalent to $B^{\rho,0}_{k}\otimes_{k}(B^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$.
@@
    $(B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}$ is derived equivalent to $(B^{\rho,0}_{k}\otimes_{k}(B^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}.$
@@
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then
@@
    $B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$ is splendidly Rickard equivalent to $B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$.
@@
    $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$ is splendidly Rickard equivalent to $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$.
See Definitions 4.9 and 4.11 below for splendid Rickard equivalences between general interior $P$-algebras. In [21], Kleshchev and Livesey defined the notion of a $d$-Rouquier $\bar{p}$-core for $d\in\mathbb{N}$ (see 5.16 below) and proved the following result:
Theorem 1.6 ([21, Theorem D], [22, Theorem A]).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\rho,d}_{R}$ is Morita superequivalent to $B^{\rho,0}_{R}\otimes_{R}(B^{\emptyset,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$ via a bimodule with endopermutation source.
For the definition of a Morita superequivalence, see [22, §3.3] or 2.4 below. Following [22] we call the block $e_{\rho,d}$ of $R\widetilde{S}_{n}$ in Theorem 1.6 an RoCK block of $\widetilde{S}_{n}$. Since $e_{\rho,d}$ is also a block idempotent of $R\widetilde{A}_{n}$, following [22] we also called the the block $e_{\rho,d}$ of $R\widetilde{A}_{n}$ an RoCK block of $\widetilde{A}_{n}$. Using Theorem 1.6, Kleshchev and Livesey proved Broué’s abelian defect group conjecture for RoCK blocks of $R\widetilde{S}_{n}$ and $R\widetilde{A}_{n}$:
Theorem 1.7 ([22, Theorem B]).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\rho,d}_{R}$ and $(B^{\rho,d}_{R})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\rho,d}_{R}$ and $(b^{\rho,d}_{R})_{\bar{0}}$ respectively.
In this paper we prove that in most cases Theorems 1.6 and 1.7 descend to $R_{p}$:
@@
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then $B^{\rho,d}_{R_{p}}$ is Morita superequivalent to $B^{\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\emptyset,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})$.
Corollary 7.5 (a corollary of Theorem 1.8) answers a question of Kleshchev and Livesey proposed in [22, Remark 8.15] in the highest generality of their question.

### v5

@@ Abstract
Recently, Kleshchev and Livesey [22, 23] proved the existence of RoCK $p$-blocks for double covers of symmetric and alternating groups over large enough coefficient rings. They proved that these RoCK blocks of double covers are Morita equivalent to standard “local” blocks via bimodules with endopermutation source. Based on this, Kleshchev and Livesey proved that RoCK blocks are splendidly Rickard equivalent to their Brauer correspondents. The analogous result for blocks of symmetric groups, a theorem of Chuang and Kessar, was an important step in Chuang and Rouquier ultimately proving Broué’s abelian defect group conjecture for symmetric groups. In this paper we show that the Morita and splendid Rickard equivalences constructed by Kleshchev and Livesey descend to the ring $\mathbb{Z}_{p}$ of $p$-adic integers, hence prove Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture for these RoCK blocks.
@@ 1 Introduction
Comparing to the refined version, the usual Broué conjecture is with the assumption that the complete discrete valuation ring is large enough for $G$. Hence Conjecture 1.1 contains more local-global information and has more applications; see [20, Corollary 1.9], [1, Theorem 1.4] and [15, Theorem 1]. Conjecture 1.1 has been proved for blocks with cyclic defect groups in [20], blocks with a Klein four defect group in [13], blocks of alternating groups in [14], blocks of ${\rm SL}_{n}(q)$ and ${\rm GL}_{n}(q)$ in defining characteristic (see [16]), and unipotent blocks of ${\rm GL}_{n}(q)$ in non-defining characteristics (see [17]). Conjecture 1.1 has also been verified for certain blocks of $p$-nilpotent and $p$-solvable groups in [3], [27] and [28]. Since any field is a splitting field for symmetric groups, Conjecture 1.1 - which in the symmetric group case is equivalent to the usual Broué conjecture - has been proved by Chuang and Rouquier [7, Theorem 7.6].
@@
For a non-negative integer $n$, denote by $S_{n}$ the symmetric group on $n$ letters and by $\widetilde{S}^{\eta}_{n}$ ($\eta\in\{\pm\}$) one of the two double covers $\widetilde{S}_{n}^{+}$ and $\widetilde{S}_{n}^{-}$ of $S_{n}$; see 2.13. Denote by $A_{n}$ the corresponding alternating group and by $\widetilde{A}_{n}$ its double cover. In this paper we provide additional evidence for Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture by considering blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$.
@@
Let $\mathcal{O}$ be a complete discrete valuation ring. Denote by $z$ the central element in $\widetilde{S}^{\eta}_{n}$ of order $2$. Since $\mathcal{O}\widetilde{S}^{\eta}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}S_{n}$ and $\mathcal{O}\widetilde{A}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}A_{n}$, Conjecture 1.1 is true for blocks of $\mathcal{O}\widetilde{S}^{\eta}_{n}(1+z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1+z)/2$. So we restrict our attention in this paper to blocks of $\mathcal{O}\widetilde{S}^{\eta}_{n}(1-z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1-z)/2$, which are called the spin blocks of the symmetric and alternating groups, respectively. For block with trivial defect group, Conjecture 1.1 is trivially true, hence we only need to consider Conjecture 1.1 for block with nontrivial defect groups. Let $K$ denote the field of fractions of $\mathcal{O}$ and let $k$ denote the residue field of $\mathcal{O}$, i.e. $k=\mathcal{O}/J(\mathcal{O})$. Let $p:={\rm char}(k)$, a prime number. For the rest of this introduction we assume that $K$ has characteristic 0, and when dealing with a finite group $G$, we will assume that $K$ (and consequently $\mathcal{O}$) contains a primitive $|G|$-th root of unity; following [22, 23] we also assume that 1-1 and $2$ have square roots (say $\sqrt{-1}$ and $\sqrt{2}$) in $K$ (and hence $\mathcal{O}$). Let $\mathbb{Z}_{p}$ denote the ring of $p$-adic integers, $\mathbb{Q}_{p}$ the field of $p$-adic numbers and $\mathbb{F}_{p}$ the field with $p$ elements. Since $z$ is a central element of order $2$, if $p=2$, then $\widetilde{S}^{\eta}_{n}$ (resp. $\widetilde{A}_{n}$) satisfies Conjecture 1.1 if and only if $S_{n}\cong\widetilde{S}^{\eta}_{n}/\langle z\rangle$ (resp. $A_{n}\cong\widetilde{A}_{n}/\langle z\rangle$) satisfies Conjecture 1.1; see [31, Lemma 10.2.11] (we note that the blanket assumption in [31] that the coefficient ring is large enough is not used in the proof of [31, Lemma 10.2.11]). Hence Conjecture 1.1 holds for $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ if $p=2$. So from now on we assume that $p\geq 3$.
@@
Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$ and let $e_{z}:=(1-z)/2\in R\widetilde{S}^{\eta}_{n}$. From [22, 23] we see that it is important to consider $R\widetilde{S}^{\eta}_{n}$ as a superalgebra via
$(R\widetilde{S}^{\eta}_{n})_{\bar{0}}:=R\widetilde{A}_{n}~~~{\rm and}~~~(R\widetilde{S}^{\eta}_{n})_{\bar{1}}:=R(\widetilde{S}^{\eta}_{n}\backslash\widetilde{A}_{n}).$
@@
Recall from [23] that the spin superblock algebras (i.e. the indecomposable superideals) of $R\widetilde{S}^{\eta}_{n}e_{z}$ are labeled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core ($p$-bar core) and $d\in\mathbb{N}$ (called the weight of the block) such that $n=|\rho|+dp$. Denote such a superblock algebra by $B^{\eta,\rho,d}_{R}$ and denote by $e_{\rho,d}$ the central idempotent of $R\widetilde{S}^{\eta}_{n}$ such that $B^{\eta,\rho,d}_{R}=R\widetilde{S}^{\eta}_{n}e_{\rho,d}$. We prove (in Proposition 6.8 below) that the idempotent $e_{\rho,d}$ is contained in the subring $R_{p}\widetilde{S}^{\eta}_{n}$. Hence we can define $B^{\eta,\rho,d}_{R_{p}}:=R_{p}\widetilde{S}^{\eta}_{n}e_{\rho,d}$. Note that the case $d=0$ corresponds to the defect zero situations, so we assume that $d>0$ in the rest of this section. In this case $B^{\eta,\rho,d}_{R}$ is in fact a spin block algebra (and not just a superblock algebra) of $R\widetilde{S}^{\eta}_{n}$.
@@
We denote by $D^{\rho,d}$ a defect group of $B^{\eta,\rho,d}_{R}$. The defect group $D^{\rho,d}$ of $B^{\eta,\rho,d}_{R}$ is abelian if and only if $d<p$. Moreover, if $d<p$, then $D^{\rho,d}\cong(C_{p})^{\times d}$. By [23, Remark 4.4], the idempotent $e_{\rho,d}$ is contained in the subalgebra $R\widetilde{A}_{n}e_{z}$, and hence (by Proposition 6.8) in $R_{p}\widetilde{A}_{n}e_{z}$. By [23, Theorem 4.5] (or [19, Propsition 3.16]), as an element of $R\widetilde{A}_{n}$, $e_{\rho,d}$ is a single block unless $d=0$ and $\rho$ is even. So $B^{\eta,\rho,d}_{R}$ can be regarded as an $R$-superalgebra with even part $(B^{\eta,\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}$. We denote $R_{p}\widetilde{A}_{n}e_{\rho,d}$ by $(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$. It is known that $D^{\rho,d}$ is also a defect group of $B^{\eta,\rho,d}_{\bar{0}}$ (see [19, Proposition 3.16] or Theorem 6.7 below). Let $B^{\eta,\rho,d}_{R}$ (resp. $(b^{\eta,\rho,d}_{R})_{\bar{0}}$) be the Brauer correspondent of $B^{\eta,\rho,d}_{R}$ (resp. $(B^{\eta,\rho,d}_{R})_{\bar{0}}$) in $RN_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})$ (resp. $RN_{\widetilde{A}_{n}}(D^{\rho,d})$). Similarly, we let $b^{\eta,\rho,d}_{R_{p}}$ (resp. $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) be the Brauer correspondent of $B^{\eta,\rho,d}_{R_{p}}$ (resp. $(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) in $R_{p}N_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})$ (resp. $R_{p}N_{\widetilde{A}_{n}}(D^{\rho,d})$). Note that $(b^{\eta,\rho,d}_{R})_{\bar{0}}$ (resp. $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) is indeed the even part of the superalgebra $b^{\eta,\rho,d}_{R}$ (resp. $b^{\eta,\rho,d}_{R_{p}}$). If $d>0$, we have $b^{\eta,\rho,d}_{R}=RN_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})e_{\rho,0}$, $(b^{\eta,\rho,d}_{R})_{\bar{0}}=RN_{\widetilde{A}_{n}}(D^{\rho,d})e_{\rho,0}$, $b^{\eta,\rho,d}_{R_{p}}=R_{p}N_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})e_{\rho,0}$ and $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}=R_{p}N_{\widetilde{A}_{n}}(D^{\rho,d})e_{\rho,0}$; see Theorem 6.5 (ii) below.
Let $\mathcal{T}_{n,R}$ be the twisted group superalgebra defined by Kleshchev and Livesey [22, 23]; see 2.7 below. For a superalgebra $A$, the twisted wreath superproduct $A\wr_{s}\mathcal{T}_{n,R}$ is defined in [23, §3.7] (see also 2.7 below). In order to consider the refined Broué conjecture, for $\eta\in\{\pm\}$, we define (in 3.1) the $\eta$-twisted group superalgebra $\mathcal{T}^{\eta}_{n,R}$ and the $\eta$-twisted wreath superproduct $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$. We prove in Proposition 3.2 that $\mathcal{T}_{n,R}\cong\mathcal{T}^{\eta}_{n,R}\cong R\otimes_{R_{p}}\mathcal{T}^{\eta}_{n,R_{p}}$ and $A\wr_{s}\mathcal{T}_{n,R}\cong A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ as $R$-superalgebras. Note that if $\sqrt{-1}\notin R_{p}$, then it seems that $\mathcal{T}^{+}_{n,R_{p}}$ and $\mathcal{T}^{-}_{n,R_{p}}$ are not isomorphic as $R$-algebras.
In [22], Kleshchev and Livesey proved the following structure theorem of $b^{\eta,\rho,d}_{k}$:
Theorem 1.2 ([22, Theorem A]).
@@
    $b^{\eta,\rho,d}_{k}\cong B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})$ as $k$-superalgebras.
@@
    $(b^{\eta,\rho,d}_{k})_{\bar{0}}\cong(B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}))_{\bar{0}}$ as $k$-algebras.
Based on the above theorem we prove a similar structure theorem of $b^{\eta,\rho,d}_{\mathbb{F}_{p}}$:
@@
    $b^{\eta,\rho,d}_{\mathbb{F}_{p}}\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})$ as $\mathbb{F}_{p}$-superalgebras.
@@
    $(b^{\eta,\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}\cong(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}))_{\bar{0}}$ as $\mathbb{F}_{p}$-algebras.
In [22], Kleshchev and Livesey proved the following analogue of “the Rickard and Mǎrcuş step for symmetric groups” for double covers of symmetric and alternating groups:
Theorem 1.4 ([22, Theorem C]).
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
  1. 1.
    $B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})$ is derived equivalent to $B^{\eta,\rho,0}_{k}\otimes_{k}(B^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})$.
  2. 2.
    $(B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}))_{\bar{0}}$ is derived equivalent to $(B^{\eta,\rho,0}_{k}\otimes_{k}(B^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}))_{\bar{0}}.$
@@
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
@@
    $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})$ is splendidly Rickard equivalent to $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})$.
@@
    $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}))_{\bar{0}}$ is splendidly Rickard equivalent to $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}))_{\bar{0}}$.
See Definitions 5.8 and 5.13 below for splendid Rickard equivalences between general interior $P$-algebras. In [22], Kleshchev and Livesey defined the notion of a $d$-Rouquier $\bar{p}$-core for $d\in\mathbb{N}$ (see 6.16 below) and proved the following result:
Theorem 1.6 ([22, Theorem D], [23, Theorem A and Lemma 9.2]).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\eta,\rho,d}_{R}$ is Morita superequivalent to $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d,R})$ via a bimodule $X^{R}$ with an endopermutation source $U^{R}$.
For the definition of a Morita superequivalence, see [23, §3.3] or 2.4 below. Following [23] we call the block $e_{\rho,d}$ of $R\widetilde{S}^{\eta}_{n}$ in Theorem 1.6 an RoCK block of $\widetilde{S}^{\eta}_{n}$. Since $e_{\rho,d}$ is also a block idempotent of $R\widetilde{A}_{n}$, following [23] we also called the the block $e_{\rho,d}$ of $R\widetilde{A}_{n}$ an RoCK block of $\widetilde{A}_{n}$. Using Theorem 1.6, Kleshchev and Livesey proved Broué’s abelian defect group conjecture for RoCK blocks of $R\widetilde{S}^{\eta}_{n}$ and $R\widetilde{A}_{n}$:
Theorem 1.7 ([23, Theorem B]).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\eta,\rho,d}_{R}$ and $(B^{\eta,\rho,d}_{R})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\eta,\rho,d}_{R}$ and $(b^{\eta,\rho,d}_{R})_{\bar{0}}$ respectively.
In this paper we prove that Theorems 1.6 and 1.7 descend to $R_{p}$:
@@
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\eta,\rho,d}_{R_{p}}$ is Morita superequivalent to $B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})$ via a bimodule $X^{R_{p}}$ with an endopermutation source $U^{R_{p}}$, such that $R\otimes_{R_{p}}X^{R_{p}}$ and $R\otimes_{R_{p}}U^{R_{p}}$ are isomorphic to the $X^{R}$ and $U^{R}$ in Theorem 1.6, respectively.
Corollary 8.5 (a corollary of Theorem 1.8) answers a question of Kleshchev and Livesey proposed in [23, Remark 8.15] in the highest generality of their question.